Questions 3 and 4 are based on the following information: The demand and supply functions for free-range chickens are Pd = 50 − 0,6Q and Ps = 20 + 0,4Q, where P is the price and Q the quantity. Question 3 The equilibrium price and quantity are [1] P = 8,00; Q = 70. [2] P = 32,00; Q = 30. [3] P = 48,50; Q = 3. [4] P = 68,00; Q = 30. 10 DSC1520/001/3 Question 4 If free-range chickens are subsidised by R4, the equilibrium price and quantity are [1] P = 17,20; Q = 3. [2] P = 29,60; Q = 34. [3] P = 34,40; Q = 26. [4] P = 84,00; Q = 170


1
Expert's answer
2021-05-25T09:00:53-0400

Answer 3 :

To find equilibrium price and quantity,

put "P_d=P_s"

"\Rightarrow 50-0,6Q=20+0,4Q\\\Rightarrow Q=30\\\therefore P=50-(0,6)(30)\\\Rightarrow P=50-18,00=32,00"


Answer: "\boxed{[2] P=32,00\ \ ;\ Q=30}"



Answer 4:

Free range chicken are subsidised by

"\Rightarrow \text{New supply function}=20+0,4Q-4"

"= 16+0,4Q"

"\therefore" For equilibrium , put

"\Rightarrow 16+0,4Q=50-0,6Q\\\Rightarrow Q= 34\\\ \\\therefore16+0,4Q=16+0,4(Q=34)=29,60"


Answer: "\boxed{[2]P=29,60\ \ ;\ Q=34}"